### v1

# 1bit-Merging: Dynamic Quantized Merging for Large Language Models

Abstract
Recent advances in large language models have led to specialized models excelling in specific domains, creating a need for efficient model merging techniques. While traditional merging approaches combine parameters into a single static model, they often compromise task-specific performance. However, task-specific routing methods maintain accuracy but introduce substantial storage overhead. We present 1bit-Merging, a novel framework that integrates task-specific routing with 1-bit quantized task vectors to balance performance and storage efficiency. Our approach leverages the observation that different task-specific models store knowledge in distinct layers—chat models primarily in attention layers and math/code models in MLP layers—enabling targeted compression strategies. Through extensive experiments with LLaMA2 and Mistral model families across chat, mathematical reasoning, and code generation tasks, we demonstrate that 1bit-Merging achieves comparable or superior performance to existing methods while significantly reducing storage requirements. Our framework offers a practical solution for combining specialized models while maintaining their individual strengths and addressing the storage challenges of current approaches.
1bit-Merging: Dynamic Quantized Merging for Large Language Models
Shuqi Liu1,2, Han Wu2,, Bowei He1, Zehua Liu2, Xiongwei Han2, Mingxuan Yuan2, Linqi Song1, 1 Department of Computer Science, City University of Hong Kong 2 Huawei Noah’s Ark Lab {shuqiliu4-c, boweihe2-c}@my.cityu.edu.hk wu.han1@huawei.com linqi.song@cityu.edu.hk
footnotetext: Corresponding author.
1 Introduction
Figure 1: While individually fine-tuned models excel only in their specialized domains, our 1bit-Merging achieves superior performance across all domains.
Large language models have achieved remarkable progress, demonstrating strong performance on a wide range of tasks Touvron et al. (2023); Zhao et al. (2023). As researchers continue to fine-tune these models for specific domains, there is a growing need to combine their specialized capabilities into a single model (Yang et al., 2024; Goddard et al., 2024). While multi-task learning offers one solution Sanh et al. (2022); Fifty et al. (2021), it requires extensive computational resources and simultaneous access to all task-specific datasets. Recent advances in parameter-space model merging Wortsman et al. (2022); Ilharco et al. (2023); Yadav et al. (2023); Yu et al. (2024b) provide an efficient alternative - by directly operating on model parameters, these methods preserve data privacy and eliminate the need for expensive retraining.
Traditional model merging approaches Ilharco et al. (2023); Yadav et al. (2023); Yu et al. (2024b) typically combine the parameters of multiple fine-tuned models, or expert models, into a single static model without additional training, thereby enabling efficient multi-task functionality. However, merging models from different domains often sacrifices task-specific performance, resulting in a noticeable gap compared to individual expert models. In contrast, merging with task-specific routing (Muqeeth et al., 2024; Lu et al., 2024) dynamically prioritizes relevant task vectors based on input data, effectively maintaining accuracy by isolating task-specific parameters. However, this routing-based merging strategy introduces substantial storage overhead, as it necessitates the preservation of all task vectors to ensure task relevance and performance. Thus, despite their ability to uphold model accuracy, task-specific routing methods face severe storage challenges, limiting their scalability and practicality in resource-constrained environments.
To effectively balance performance and storage efficiency, we introduce 1bit-Merging, a novel dynamic merging framework that integrates task-specific routing with 1-bit quantized task vectors. Recognizing the substantial redundancy inherent within task vectors, we implement 1-bit compression, which significantly reduces storage requirements without notably compromising the model’s effectiveness. Notably, we observe that different task-specific models store knowledge in distinct layers. For chat-oriented models, knowledge is predominantly stored in attention layers, enabling us to compress MLP layers. Conversely, for math and code-related models, knowledge is primarily stored in MLP layers, allowing us to compress attention layers. Building upon the compressed task vectors, our framework employs task-specific routing to establish a task-specific base model. This base model serves as the foundation for integrating the remaining compressed task vectors, ensuring that each task leverages the most relevant and efficient parameters. 1bit-Merging thus offers a balanced solution that maintains the performance advantages of task-specific routing while addressing the storage inefficiencies of existing approaches.
To empirically demonstrate the effectiveness of 1bit-Merging, we conduct extensive experiments by combining it with existing model merging approaches. We merged three widely adopted fine-tuned models—specializing in general knowledge (Chat), mathematical reasoning (Math), and code generation (Code)—derived from the LLaMA2-7B/13B and Mistral 7B families. Through extensive experiments across multiple tasks, we demonstrate that our approach not only outperforms traditional model merging methods but also achieves better storage efficiency than task-specific routing approaches. Our task vector compression method effectively preserves and often enhances the capabilities of fine-tuned models. By integrating these compressed task vectors with our dynamic routing strategy, our 1bit-Merging method achieves superior performance. As illustrated in Figure 1, compared to individually fine-tuned models, our 1bit-Merging method delivers better performance across all domains.
To sum up, our contributions include: (1) We propose a novel dynamic merging framework that integrates task-specific routing with 1-bit quantized task vectors. (2) We empirically demonstrate that different task-specific models store knowledge in distinct layers, enabling targeted compression strategies based on specific compression position. (3) Through comprehensive evaluations, we validate that our proposed method enhances model merging performance across various domains and achieves better storage efficiency than task-specific routing approaches.
2 Related Work
Modeling merging (Yang et al., 2024; Goddard et al., 2024) has attracted much attention for its several advantages: 1) it significantly reduces the storage and deployment expenses by consolidating multiple models into a single one (Ilharco et al., 2023; Yu et al., 2024a); 2) it operates in a plug-and-play manner, eliminating the need for additional training(Wortsman et al., 2022; Yu et al., 2024b); and 3) it can improve performance on individual tasks(Wortsman et al., 2022; Ilharco et al., 2023; Yadav et al., 2023) while also improving out-of-domain generalization(Cha et al., 2021; Jin et al., 2023). Based on whether the merged model remains consistent across all samples or tasks, model merging approaches can be categorized into two types: static model merging and dynamic model merging.
2.1 Static Model Merging
Static model merging primarily explores general strategies for combining models, such as Average Merging (Wortsman et al., 2022), Task Arithmetic (Ilharco et al., 2023), and Ties-Merging (Yadav et al., 2023). Notably, Wortsman et al. (2022) first demonstrated that even a straightforward weight averaging of base models can improve both the performance and robustness of downstream tasks. Building on this, Task Arithmetic (Ilharco et al., 2023) refines the merging process by introducing task vectors, proposing that simple arithmetic operations on these vectors can effectively modify models and yield a better merged model. Expanding upon the concept of task vectors, methods like DARE (Yu et al., 2024b) and Ties (Yadav et al., 2023) adopt pruning-then-scaling techniques to merge task vectors, based on the premise that not all parameters equally contribute to the final performance. However, static merging of models from different domains often sacrifices task-specific performance to strike a balance between generalization capacity and task-specific effectiveness. To this end, we explore the dynamic model merging in this work.
2.2 Dynamic Model Merging
Additionally, another line of studies focuses on routing-based model merging. For instance, SMEAR (Muqeeth et al., 2024) propose a routing-based merging paradigm where parameter fusion is implemented through weighted averaging guided by router input distributions across expert modules, maintaining computational efficiency comparable to singular expert operations. Extending this paradigm, Twin-Merging (Lu et al., 2024) develop an adaptive knowledge integration framework that dynamically reconciles task-shared and task-specific representations via routing mechanisms during inference. (Tang et al., 2024) advance this domain through a Transformer-based dynamic composition architecture, with empirical analysis revealing disproportionate parameter modification magnitudes between linear and nonlinear layers during fine-tuning - a critical factor affecting integration efficacy. Notwithstanding their methodological advancements, extant routing-based merging frameworks incur significant storage demands for task vector retention, which motivates the development of our computationally efficient model merging framework in this work.
General Knowledge Mathmetical Reasoning Code Generation
Models Position MMLU HellaSwag TruthfulQA GSM8K MATH MBPP HumanEval Average
/ 46.38 57.79 45.17 23.43 4.86 0.3 0.6 25.50
Attention 42.12 55.30 41.13 22.74 5.08 0.0 0.0 23.70
MLP 46.40 58.23 43.70 21.15 4.94 0.0 0.0 25.30
Chat Linear 45.61 58.38 42.96 20.70 4.76 0.0 0.0 24.63
/ 40.05 56.30 32.56 48.60 8.50 21.8 12.8 31.52
Attention 41.03 56.68 34.52 47.46 8.42 23.1 14.0 32.17
MLP 41.85 57.68 33.17 44.88 7.06 25.3 14.0 31.99
Math Linear 42.76 57.96 34.52 42.29 6.84 24.3 15.2 31.98
/ 40.76 57.87 33.17 7.13 3.62 26.8 5.5 24.98
Attention 41.36 57.81 32.80 8.19 3.36 27.3 11.0 25.97
MLP 41.08 57.80 33.29 5.23 3.16 14.0 2.4 22.42
Code Linear 41.50 57.60 33.54 7.51 3.06 0.0 0.0 20.46
Table 1: Impact of different layer types on compression performance across Math, Code, and Chat expert models derived from LLaMA-2 7B model. Attention layer compression performs best for Math and Code models, while MLP layer compression yields optimal results for Chat model.
3 Method
3.1 Traditional Model Merging and Merging with Task-Specific Routing
We first compare traditional model merging with the merging strategies combined with task-specific routing. Starting with $K$ fine-tuned models $\{\theta^{t_{1}}_{\text{SFT}},\theta^{t_{2}}_{\text{SFT}},\ldots,\theta^{t_{K}}_{\text{SFT}}\}$ derived from a common pre-trained backbone $\theta_{\text{PRE}}$, each task vector is defined as the difference between and after finetuning:
$\delta_{t_{k}}=\theta^{t_{k}}_{\text{SFT}}-\theta_{\text{PRE}},\quad\text{for }k\in\{1,\ldots,K\}.$
Traditional merging aggregates all task vectors to construct a single static merged model:
$\theta_{\text{merged}}^{\text{traditional}}=\theta_{\text{PRE}}+\sum_{k=1}^{K}\delta_{t_{k}}.$
In contrast, merging with routing leverages task specialized knowledge tailored to each input $\mathbf{x}$ through a router:
$k^{*}=\text{Router}(\mathbf{x}),\quad\theta_{\text{merged}}^{\text{routing}}=\theta_{\text{PRE}}+\delta_{t_{k^{*}}}.$
As shown in Figure 2(a), traditional merging methods like Task Arithmetic Ilharco et al. (2023) are highly impacted by task interference: merging dissimilar tasks (math and chat) degrades performance more than similar tasks (math and code), and adding more tasks exacerbates the decline. In contrast, task-specific routing prevents interference and preserves the accuracy of fine-tuned models.
(a) Gaps from merging with task-specific routing.
(b) Impact of merging with the fine-tuned model as base.
Figure 2: Performance of merged model in traditional model merging and merging with task-specific settings.
3.2 Impact of Base Model in Model Merging
Building on the strength of task-specific routing, we further investigate the impact of selecting different base models in model merging. Specifically, we compare the performance outcomes when employing a pre-trained backbone $\theta_{\text{PRE}}$ versus a math-finetuned model $\theta_{\text{Math}}=\theta_{\text{PRE}}+\delta_{t_{\text{Math}}}$ as the base model for mathematical input data. As shown in Figure 2(b), substituting the pre-trained backbone with a math-finetuned model significantly enhances performance on the GSM8K Cobbe et al. (2021) dataset. Consequently, initializing the base model with task-specific fine-tuning allows the merged model to more effectively solve the corresponding task. Despite surpassing individually fine-tuned models, utilizing a task-specific base model concurrently increases storage requirements due to the necessity of loading multiple task vectors.
3.3 1bit-Merging
To balance performance with storage efficiency, we introduce 1bit-Merging, a dynamic merging method that combines task-specific routing with 1-bit quantized task vectors.
Compression of Task Vectors.
Recognizing the substantial redundancy within task vectors, we apply 1-bit weight quantization to the task vectors, converting the weight matrices in Linear layers from FP32/16 precision to a 1-bit format. In this quantization process, each element of the task vector is set to either +1 or -1. To maintain performance despite the aggressive compression, we scale the binary weight matrices with scalar values in FP16 format. This scaling ensures that the quantized weights preserve the original weight $L_{2}$ norm, thereby maintaining model performance after the extremely low-bit compression Liu et al. (2024). The scaling factor $\alpha$ is computed as:
$\alpha=\frac{\|\mathbf{W}\|_{1}}{{m\cdot n}}$
where $m$ and $n$ are the dimensions of the weight matrix. Using this scaling factor, the transformed task vector $\tilde{\delta}_{t_{k}}$ is defined as:
$\tilde{\delta}_{t_{k}}=\alpha*\text{Sign}(\delta_{t_{k}})$ (1)
Compression Position.
We examine 1-bit compression effectiveness across different model components by selectively compressing Attention layers, MLP layers, and all Linear layers. As shown in Table 1, each model type exhibits distinct compression characteristics. Math and Code models achieve optimal performance when only attention layers are compressed while Chat models perform best with MLP layer compression. Remarkably, the compressed versions of both Math and Code models demonstrate slight performance improvements over their original fine-tuned counterparts These patterns reveal a fundamental difference in how knowledge is distributed within the model architecture: task-specific capabilities (like mathematical reasoning and code generation) primarily reside in MLP layers, while attention layers appear to store more general, transferable knowledge. Based on these insights, we strategically apply attention layer quantization to Math and Code models, and MLP layer quantization to Chat models when compressing task vectors.
Dynamic Routing and Merging.
Our merging framework then incorporates the compressed task vectors with a task-specific routing mechanism, where a trained router analyzes the input data $\mathbf{x}$ to produce a probability distribution $\mathbf{p}$ across different tasks. The most relevant task vector $\delta_{t_{k}^{*}}$ is selected based on the highest probability $k^{*}$ and added to the pre-trained model parameters $\theta_{\text{PRE}}$ to form a task-specific base model $\theta_{\text{base}}=\theta_{\text{PRE}}+\delta_{t_{k}^{*}}$. Finally, we apply Ties Merging Yadav et al. (2023) to integrate the remaining compressed task vectors into $\theta_{\text{base}}$, resulting in a comprehensive model that dynamically adapts to the most relevant tasks while maintaining overall performance through efficient parameter integration.
4 Experiments
4.1 Experimental Setup
Baselines.
We evaluate the effectiveness of our 1bit-Merging method through a two-step process. First, we assess the performance of compressed task vectors by comparing compressed expert models— which integrate these vectors with the pretrained backbone—against the individually fine-tuned models. Second, we benchmark 1bit-Merging against existing model merging methods: Task Arithmetic, Ties-Merging, and DARE. Task Arithmetic (Ilharco et al., 2023) enhances the merging process by introducing task vectors, suggesting that simple arithmetic operations on these vectors can effectively merge models. Building on the concept of task vectors, both DARE (Yu et al., 2024b) and Ties-Merging (Yadav et al., 2023) employ pruning-then-scaling methods to merge task vectors, based on the assumption that not all parameters contribute equally to the final performance.
Method General Knowledge Mathmetical Reasoning Code Generation Average
MMLU HellaSwag TruthfulQA GSM8K MATH MBPP HumanEval
Finetuned Models
Chat 46.38 57.79 45.17 23.43 4.86 0.3 0.6 25.50
Math 40.05 56.30 32.56 48.60 8.50 21.8 12.8 31.52
Code 40.76 57.87 33.17 7.13 3.62 26.8 5.5 24.98
Compressed Expert Models
1bit-Chat 46.40 58.23 43.70 22.74 5.08 0.0 0.0 25.55
1bit-Math 41.03 56.68 34.52 47.46 8.42 23.1 14.0 32.17
1bit-Code 41.36 57.81 32.80 8.19 3.36 27.3 11.0 25.97
Model Merging Methods
Task Arithmetic 41.50 49.63 37.45 47.34 6.46 13.5 7.3 29.03
Ties-Merging 45.75 56.63 40.02 46.93 7.74 29.1 17.1 34.75
DARE 46.81 57.57 38.19 44.05 6.98 31.6 18.9 34.87
1bit-Merging 47.23 58.04 45.37 48.52 9.04 30.1 18.9 36.74
Table 2: Performance evaluation of merged LLaMA2-7B Models (Chat, Math, Code) across 7 task-specific datasets
Method General Knowledge Mathmetical Reasoning Code Generation Average
MMLU HellaSwag TruthfulQA GSM8K MATH MBPP HumanEval
Finetuned Models
Chat 59.05 65.97 55.69 42.53 9.16 49.6 42.7 46.37
Math 60.77 58.68 44.68 63.38 22.74 38.1 23.8 44.59
Code 50.58 53.19 45.29 31.69 4.84 50.9 40.9 39.63
Compressed Expert Models
1bit-Chat 54.71 58.00 53.24 38.51 8.14 47.6 35.4 42.23
1bit-Math 60.11 58.48 41.13 58.76 19.92 41.1 25.6 43.59
1bit-Code 47.88 53.51 42.96 27.29 5.16 41.6 33.5 35.99
Model Merging Methods
Task Arithmetic 47.34 46.80 41.00 52.16 13.26 32.1 29.9 39.17
Ties-Merging 57.20 57.59 48.71 55.50 15.00 48.4 41.5 46.30
DARE 55.36 55.77 42.84 57.39 15.00 49.4 39.0 44.97
1bit-Merging 56.37 57.87 53.06 60.42 20.60 48.6 42.1 48.43
Table 3: Performance evaluation of merged Mistral 7B Models (Chat, Math, Code) across 7 task-specific datasets
Benchmark.
Our experimental evaluation focuses on three model families: LLaMA-2 7B Touvron et al. (2023), Mistral 7B Jiang et al. (2023), and LLaMA-2 13B Touvron et al. (2023), each covering distinct specializations in: general knowledge (Chat), mathematical reasoning (Math), and code generation (Code). We assess performance using seven benchmark datasets across three domains: MMLU Hendrycks et al. (2020), HellaSwag Zellers et al. (2019) and TruthfulQA Lin et al. (2022) for assessing general knowledge and reasoning capabilities; GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021) for testing mathematical reasoning proficiency; and HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) for evaluating code generation abilities. To ensure consistent and unbiased assessment, model performance is evaluated using zero-shot accuracy, with pass@1 rate specifically measuring code generation correctness.
4.2 Main Results
Using the Chat111huggingface.co/meta-llama/Llama-2-7b-chat-hf, Math222huggingface.co/TIGER-Lab/MAmmoTH-7B, and Code333huggingface.co/mrm8488/llama-2-coder-7b fine-tuned models derived from the same LLaMA2-7B pretrained model444huggingface.co/meta-llama/Llama-2-7b-hf, we first compare the performance of the compressed expert models with the individually fine-tuned models to demonstrate the effectiveness of our task vector compression. Subsequently, we evaluate the performance of our 1bit-Merging model against existing task-vector-based model merging methods. Table 3 presents a comprehensive comparison across seven datasets.
Method General Knowledge Mathmetical Reasoning Code Generation Average
MMLU HellaSwag TruthfulQA GSM8K MATH MBPP HumanEval
Finetuned Models
Chat 53.17 60.73 40.88 32.37 6.70 16.5 7.9 31.18
Math 52.73 61.10 37.09 55.50 10.84 28.8 15.9 37.42
Code 52.65 60.42 40.64 27.29 5.74 21.3 10.4 31.21
Compressed Expert Models
1bit-Chat 53.47 60.40 41.25 31.39 6.58 28.1 5.5 32.38
1bit-Math 53.52 61.39 35.86 56.71 10.38 31.6 16.5 37.80
1bit-Code 53.06 60.08 39.17 26.61 6.12 21.6 12.2 31.26
Model Merging Methods
Task Arithmetic 52.22 57.52 41.49 49.89 7.32 24.1 9.1 34.95
Ties-Merging 55.48 60.65 39.05 52.46 9.90 40.4 21.3 39.89
DARE 55.65 61.66 40.51 55.19 9.08 39.1 20.1 40.18
1bit-Merging 55.48 60.22 42.21 56.65 10.40 40.6 22.0 41.07
Table 4: Performance evaluation of merged LLaMA2-13B Models (Chat, Math, Code) across 7 task-specific datasets
Compressed Expert Models.
Our task vector compression method demonstrates remarkable effectiveness, with the compressed expert models not only maintaining but often surpassing the performance of their fine-tuned counterparts. The 1bit-Chat model achieves an average score of 25.55, surpassing its fine-tuned counterpart’s 25.50. More notably, the 1bit-Math model improves from 31.52 to 32.17, while the 1bit-Code model shows the most substantial gain, increasing from 24.98 to 25.97. Moreover, the compressed expert models exhibit strong adaptability, maintaining high performance not only in their specialized domains but also across other non-specialized domains. (1) Superior performance in specialized domains: Within their specialized domains, the 1bit-Chat model enhances general knowledge capabilities with a 0.63 point improvement, the 1bit-Math model maintains near-optimal performance, and most impressively, the 1bit-Code model achieves a substantial 3.0 point boost in code generation proficiency. These results suggest that our compression method not only preserves but often enhances domain-specific expertise. (2) Superior performance in non-specialized domains: The benefits of our compression technique extend beyond primary specializations, showcasing enhanced cross-domain generalization. This is particularly evident in the compressed Math model, which achieves notable gains of 1.11 points in general knowledge and 1.25 points in code generation—domains outside its primary expertise. These consistent improvements across both specialized and general domains indicate that our compression method facilitates positive knowledge transfer.
Merging Models with 1bit-Merging.
Our 1bit-Merging method then integrates the compressed expert models through dynamic routing. As shown in Table 3, our 1bit-Merging method demonstrates superior performance compared to existing merging techniques in two significant aspects. (1) Consistent improvements across all tasks: While baseline methods exhibit distinctive domain-specific strengths - with Ties-Merging excelling in mathematical reasoning and DARE showing superior results in general knowledge and code generation - our 1bit-Merging method achieves comparable or better performance across all domains. Specifically, our 1bit-Merging method surpasses Ties-Merging by 1.45 points in mathematical reasoning and DARE by 2.71 points in general knowledge. Although we observe a modest 0.75-point decrease in code generation compared to DARE, this minor regression is primarily attributable to the lower baseline performance of the Code fine-tuned model. Notably, our method still achieves a substantial 8.35-point improvement over the original code fine-tuned model. (2) Exceeds the highest average performance of expert fine-tuned models: Perhaps most significantly, our method achieves a remarkable average performance score of 36.74 across all tasks, markedly surpassing the average best performance (35.16) of task-specific fine-tuned models—a distinctive achievement that none of the existing merging methods have accomplished. This remarkable improvement of 1.58 points over the highest individual expert scores demonstrates that our 1bit-Merging method not only preserves but actually enhances the specialized capabilities of individual models, suggesting the merged model leverages complementary strengths across different domains and maintains their peak performance.
General Knowledge Mathmetical Reasoning Code Generation
Models Position MMLU HellaSwag TruthfulQA GSM8K MATH MBPP HumanEval Average
/ 59.05 65.97 55.69 42.53 9.16 49.6 42.7 46.37
Attention 53.25 57.03 54.59 39.20 8.14 46.6 36.0 42.12
MLP 54.71 58.00 53.24 38.51 8.14 47.6 35.4 42.23
Chat Linear 54.95 58.83 53.73 38.82 8.26 47.6 37.2 42.77
/ 60.77 58.68 44.68 63.38 22.74 38.1 23.8 44.59
Attention 60.11 58.48 41.13 58.76 19.92 41.1 25.6 43.59
MLP 58.74 58.12 43.57 53.37 16.68 44.9 29.3 43.53
Math Linear 57.93 58.24 39.41 47.16 13.9 41.1 25.6 40.48
/ 50.58 53.19 45.29 31.69 4.84 50.9 40.9 39.63
Attention 47.88 53.51 42.96 27.29 5.16 41.6 33.5 35.99
MLP 45.65 54.56 42.59 34.34 6.08 47.9 28.0 37.02
Code Linear 39.99 53.80 36.47 22.52 4.82 40.6 31.1 32.76
Table 5: Impact of different layer types on compression performance on Mistral 7B models.
4.3 Using Different Model Architecture
To validate the generalizability of our method, we extend our experiments to the Mistral 7B architecture. As illustrated in Table 3, Our approach demonstrates robust performance improvements despite architectural differences. When merging Chat555huggingface.co/mistralai/Mistral-7B-Instruct-v0.1, Math666huggingface.co/TIGER-Lab/MAmmoTH2-7B, and Code777huggingface.co/Nondzu/Mistral-7B-codealpaca-lora fine-tuned models derived from pretrained backbone 888huggingface.co/mistralai/Mistral-7B-v0.1. 1bit-Merging outperforms Task Arithmetic by an average of 9.26 points, Ties-Merging by 2.13 points, and DARE by 3.46 points across all evaluated datasets. These enhancements are particularly significant in mathematical reasoning and general knowledge tasks. On the benchmark datasets MATH and GSM8K, our method respectively achieves accuracies of 20.60% and 60.42%, significantly exceeding DARE—the strongest baseline for mathematical reasoning—by 5.60 points (a 33.73% relative improvement) on MATH and by 3.03 points (a 5.28% relative improvement) on GSM8K. Additionally, on the TruthfulQA dataset, 1bit-Merging demonstrates substantial gains over all baselines (12.06 points over Task Arithmetic, 4.35 points over Ties-Merging, and 10.22 points over DARE) while performing on par with the Chat fine-tuned model. This underscores the effectiveness of 1bit-Merging in maintaining the peak performance in fine-tuned expert models. Furthermore, the average performance of our 1bit-Merging method across all datasets is 48.43, surpassing the strongest Chat fine-tuned model’s average of 46.37—a distinction not achieved by any other baseline methods.
General Knowledge Mathmetical Reasoning Code Generation
Models Position MMLU HellaSwag TruthfulQA GSM8K MATH MBPP HumanEval Average
/ 53.17 60.73 40.88 32.37 6.70 16.5 7.9 31.18
Attention 53.47 60.40 41.25 31.39 6.58 28.1 5.5 32.38
MLP 53.13 60.94 38.68 31.39 6.58 20.1 9.1 31.42
Chat Linear 53.38 61.30 39.05 31.08 6.20 32.3 4.3 32.52
/ 52.73 61.10 37.09 55.50 10.84 28.8 15.9 37.42
Attention 53.52 61.39 35.86 56.71 10.38 31.6 16.5 37.80
MLP 54.19 61.70 35.25 53.98 9.10 33.6 13.4 37.32
Math Linear 54.91 61.89 35.62 53.90 8.80 32.8 14.6 37.50
/ 52.65 60.42 40.64 27.29 5.74 21.3 10.4 31.21
Attention 53.06 60.08 39.17 26.61 6.12 21.6 12.2 31.26
MLP 53.06 60.34 38.19 22.29 4.76 5.5 12.2 28.05
Code Linear 53.42 60.60 37.09 15.84 4.58 7.8 8.5 26.83
Table 6: Impact of different layer types on compression performance on LLaMA-2 13B models.
4.4 Scaling to Larger Model Size
We further evaluate the scalability of our method using LLaMA-2 13B999huggingface.co/meta-llama/Llama-2-13b-hf architecture by merging Chat101010huggingface.co/meta-llama/Llama-2-13b-chat-hf, Math111111huggingface.co/TIGER-Lab/MAmmoTH-13B, and Code121212huggingface.co/emre/llama-2-13b-code-chat fine-tuned models. As shown in Table 4, our approach remains effective at this larger scale. The compressed expert models demonstrate consistent enhancement, achieving an average improvement of 0.54 points over their original fine-tuned counterparts. Specifically, the compressed experts show consistent improvements across all specialized domains: 1bit-Chat achieves a 0.11-point gain in general knowledge tasks, 1bit-Math demonstrates a 0.38-point improvement in mathematical reasoning, and 1bit-Code shows a notable 1.05-point enhancement in code generation. Additionally, our 1bit-Merging method demonstrates superior performance against all baseline approaches, surpassing Task Arithmetic by 6.12 points, Ties-Merging by 1.18 points, and DARE by 0.89 points on average across all evaluated datasets. When compared to DARE, the strongest baseline for mathematical reasoning and general knowledge, our method achieves notable improvements of 1.32 points (14.54% relative increase) on MATH and 1.70 points (4.20% relative increase) on TruthfulQA. In code generation tasks, our method outperforms the leading baseline, Ties-Merging, by an average of 0.45 points.
Table 7: Performance comparison between merging original and compressed fine-tuned LLaMA2-7B models.
Method General Knowledge Mathmetical Reasoning Code Generation Average
MMLU HellaSwag TruthfulQA GSM8K MATH MBPP HumanEval
LLaMA2-7B 41.50 49.63 37.45 47.34 6.46 13.5 7.3 29.03
   w/ Compressed Experts 42.49 51.45 39.17 45.72 7.70 15.8 8.5 30.09 (+1.06)
Mistral-7B 47.34 46.80 41.00 52.16 13.26 32.1 29.9 39.17
   w/ Compressed Experts 45.98 48.03 40.76 46.32 12.14 33.8 29.3 36.62 (-2.55)
LLaMA2-13B 52.22 57.52 41.49 49.89 7.32 24.1 9.1 34.95
   w/ Compressed Experts 52.57 58.01 41.49 53.60 7.48 27.3 12.2 36.09 (+1.14)
4.5 Ablation Studies
Layer Types on Compression Performance.
We further evaluate how compression position affects the performance of compressed expert models across Mistral 7B and LLaMA-2 13B architectures. As shown in Tables 5 and 6, for Chat models, while the optimal quantization locations showed slight variations—with Mistral 7B preferring all Linear layers and LLaMA2 13B favoring attention layers—the performance differences between these choices were minimal. However, for specialized Math and Code models, we find quantizing attention layers consistently to be most effective in preserving domain-specific capabilities.
Merging Compressed Expert Models.
Merging compressed expert models proves effective for Task Arithmetic but shows limitations with parameter-dropping methods like Ties-Merging and DARE. This is because parameter-dropping techniques destroy the compressed knowledge distributions essential for specialized task performance. In our experiments with LLaMA2-7B and LLaMA2-13B, where compressed experts show marginal improvements over individually fine-tuned models, merging compressed experts yields performance gains of 1.06 and 1.14 points respectively, compared to merging their fine-tuned counterparts.
Figure 3: Performance and storage trade-off of different methods in merging fine-tuned Mistral 7B models.
4.6 Performance vs. Storage Trade-offs
We demonstrate the trade-off between performance and storage requirements in Figure 3. We use Mistral 7B models, as its Code model demonstrates superior performance in code generation. Task-specific routing (Routing) achieves performance equivalent to individually fine-tuned models, serving as a performance upper bound. However, Routing requires maintaining full parameters for each task model, leading to substantial storage overhead. While model merging methods reduce storage requirements, they underperform Routing by over 4.93 points. Our 1bit-Merging strikes a favorable balance between these extremes, maintaining 94.53% of Routing performance while requiring only 66.25% of its storage. Furthermore, leveraging the observation that Chat models exhibit minimal sensitivity to compression positions, we applied quantization on all Linear layers for the Chat model. This optimization achieves an even more efficient storage reduction to 55.02% while preserving comparable performance levels.
5 Conclusion
We propose 1bit-Merging, a novel framework that effectively combines specialized language models while addressing the fundamental trade-off between performance and storage efficiency. By incorporating dynamic routing with binary quantization, our approach maintains task-specific expertise while significantly reducing storage overhead. Extensive experiments across general knowledge, mathematical reasoning, and code generation tasks demonstrate that 1bit-Merging not only preserves the specialized capabilities of individual models but often enhances their performance.
Limitations
Despite the promising results of 1bit-Merge, several limitations deserve attention. First, while our method significantly reduces storage requirements, the dynamic routing mechanism introduces computational overhead during inference, as the router must analyze each input to determine the optimal task vector and perform binary transformations. Although this overhead is relatively small, it could impact real-time applications with strict latency requirements.
Second, while compressed expert models show enhanced performance within their specialized domains, our framework may not fully capture complex cross-task interactions that could enable more sophisticated knowledge transfer, particularly for tasks requiring simultaneous expertise from multiple domains.
Third, while our 1-bit quantization approach preserves model performance for most tasks, it may not be optimal for all types of task-specific knowledge, particularly for tasks requiring high numerical precision or fine-grained reasoning.
Ethics Statement
This study utilizes publicly available datasets for our models. Prior research endeavors have generally taken ethical considerations into account. We have manually inspected a subset of samples and found no explicit ethical concerns, including violent or offensive content. Nonetheless, it is crucial to highlight that the output generated by large language models lacks the degree of control we might assume. Consequently, we are prepared to implement measures to mitigate any unforeseen outputs.